# Dynamical Tidal Response of Schwarzschild Black Holes

### v1

@@ 3.2.1 Odd sector
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]+O(\omega^{2}).$ (3.63)
@@
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j|j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.64)
@@
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=4i\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell}(\ell+1)}{\ell}\int$ ${\rm d}\tau_{1}{\rm d}\tau_{2}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}K^{(B)}_{\ell}({\omega})$ (3.65)
@@
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{4(-i)^{\ell}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)c^{[j|i_{1}]\cdots i_{\ell}}\mathcal{P}_{0j0[c}^{\text{dD}}\partial_{b]}\partial_{a}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (3.66)
@@
$\langle B_{+abc}(t,\vec{x})\rangle_{\text{in-in}}\!=\!\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell+1}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}\,$, (3.67)
and simplify the tensor structure using
$c_{[j|i_{1}]\cdots i_{\ell}}\delta^{j}_{[c}\partial_{b]}\partial_{a}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)=\partial_{a}\partial_{[b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c]|A_{\ell}}.$ (3.68)
Considering the $B_{rij}$ component of (3.66), and writing the tidal field in spherical coordinates as
@@
$\langle B_{+\,rij}(x)\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}\frac{(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}r^{2}\partial_{r}\left(r^{-2}\frac{c_{\text{ext}}}{r^{\ell+D-4}}\right).$ (3.70)
@@
$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ $=2\nabla_{r}\nabla_{[i}h_{j]0}-2\partial_{t}\nabla_{[i}h_{j]r}$ (3.71)
$=2\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\text{ext}}r^{2}\partial_{r}\big(r^{\ell-1}\big)\big(1+O(\omega r)\big)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
The $O(\omega r)$ term in $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rij}$ represents a far-zone, frequency-dependent correction to the static tidal field. It exhibits a different scaling in $r$ compared to the standard $r^{\ell}$ and $r^{-\ell-1}$ falloffs. Therefore, as long as we are concerned with matching the EFT in the intermediate zone, we can consistently neglect this contribution. Putting everything together and replacing $K_{\ell}^{(B)}$ by $\mu^{2\varepsilon}K_{\ell}^{(B)}$ (as we did in the scalar section), we obtain
@@
$\times\left(r^{\ell}+\mu^{2\varepsilon}K^{(B)}_{\ell}(\omega)\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell+3-D}\right),$
@@
where $q_{\ell}\equiv 1/\sqrt{2(\ell-1)(D-2+\ell)}$. The components (3.75) can be used to evaluate the Weyl tensor in Regge–Wheeler gauge. In particular, we find $B_{rij}$ to be
$B_{rij}^{\text{RW}}$ $=\sum_{\ell=2}^{\infty}\left[r^{2}\partial_{r}\left(r^{-2}h_{0}(t,r)\right)-\partial_{0}h_{1}(t,r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$ (3.76)
$=-\sum_{\ell=2}^{\infty}\frac{q_{\ell}\operatorname{e}^{-i\omega t}}{4}\mu^{\varepsilon}r^{1-\frac{D}{2}}\left[\left(2D-D^{2}+4r^{2}\omega^{2}\right)\Psi_{\text{RW}}(r)+4r^{2}\Psi_{\text{RW}}^{\prime\prime}(r)\right]\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$.
@@
$B_{rij}^{\text{RW}}=-\sum_{\ell=2}^{\infty}q_{\ell}\operatorname{e}^{-i\omega t}\left[(\ell-1)(\ell+2-2\varepsilon)+r^{2}\omega^{2}\right]\left(B_{\text{reg}}r^{\ell}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right)\nabla_{[i}Y_{j]}^{(T)}{}_{\ell}^{m}$. (3.77)
Since the linearized Weyl tensor is gauge invariant in flat space, we can directly compare $B_{rij}^{\text{RW}}$ with (3.72), obtained in de Donder gauge. By matching the $r^{\ell}$ and $r^{-\ell-1+2\varepsilon}$ coefficients in the two expressions, we can identify
$\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}K^{(B)}_{\ell}(\omega)=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.78)
@@
$\frac{45}{\pi}K^{(B)}_{\ell=2}(\omega)=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{787}{2520}-\log(\mu r_{s})\right).$ (3.87)
@@
$\frac{18900}{\pi}K^{(B)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}+\omega^{2}r_{s}^{9}\left(\frac{1\,727}{2\,520}-\log(\mu r_{s})\right),$ (3.88) $\frac{11113200}{\pi}K^{(B)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}+\omega^{2}r_{s}^{11}\left(\frac{237\,529}{277\,200}-\log(\mu r_{s})\right).$
@@ 3.2.2 Even sector
As in the magnetic sector, we can expand the response kernel in fourier space similarly to (3.53), where
@@
$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\tilde{\omega}(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{E}+i{\omega}r_{s}\lambda_{1,\ell}^{E}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{E}+\cdots\right],$
$\lambda_{0,\ell}^{E}$ are the static Love numbers, $\lambda_{1,\ell}^{E}$ capture the linear-in-frequency dissipative response of the object, and $\lambda_{2,\ell}^{E}$ are the dynamical Love numbers at order $O(\omega^{2})$. Since we are interested in computing $\lambda_{2,\ell}^{E}$ for Schwarzschild black holes in four-dimensional general relativity, we will set $\lambda_{0,\ell}^{E}=0$ in the following. We will allow us to consistently neglect $O(\omega^{2})$ terms in the expressions of $\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}$ and $\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle$ in (3.92).
@@
$E_{ab}$ $=C_{0a0b}=2\partial_{a}\partial_{[0}h_{b]0}-2\partial_{0}\partial_{[0}h_{b]a}=-\partial_{a}\partial_{b}h_{00}+2\partial_{0}\partial_{(a}h_{b)0}+O(\omega^{2}),$ (3.95)
$E_{A_{\ell}}$ $=-\partial_{A_{\ell}}h_{00}+2\partial_{0}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0}+O(\omega^{2}).$
@@
$E_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}}=\partial_{A_{\ell}}h_{00}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}-2\partial_{A_{\ell}}h_{00}\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{i_{\ell}0}-2\partial_{0}\partial_{A_{\ell-1}}h_{i_{\ell}0}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}+O(\omega^{2}).$ (3.96)
@@
$\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)=-\langle\partial_{a}\partial_{b}h_{+\,00}\partial_{A_{\ell}}h_{-\,00}\rangle\left(\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{00}-2\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{i_{\ell}0}\right)+O(\omega^{2}),$ (3.97)
@@
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}(t,\vec{x})=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{\rm ext}Y_{\ell}^{m}r^{\ell}=\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.99)
@@
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\frac{D-3}{D-2}\sum_{\ell=2}^{\infty}\ell!(-i)^{\ell}K_{\ell}^{(E)}(\omega)c^{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}$ (3.100)
@@
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}K_{\ell}^{(E)}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{D-3}{D-2}c_{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}.$ (3.101)
@@
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}$ $=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}+2\partial_{0}\partial_{r}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0r}+O(\omega^{2})$ (3.102)
$=-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2})=-\operatorname{e}^{-i\omega\tau}\sum_{\ell=2}^{\infty}c_{\rm ext}\,\ell(\ell-1)r^{\ell-2}Y_{\ell}^{m}+O(\omega^{2}).$

### v2

@@ 3.2.1 Odd sector
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]+O(\omega^{2}).$ (3.63)
@@
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}(t,\vec{x})=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j|j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.64)
@@
$\langle B_{+\,abc}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{4(-i)^{\ell}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)c^{[j|i_{1}]\cdots i_{\ell}}\partial_{a}\mathcal{P}_{0j0[c}^{\text{dD}}\partial_{b]}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}\,.$ (3.66)
@@
$\langle B_{+abc}(t,\vec{x})\rangle_{\text{in-in}}\!=\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{(-1)^{\ell+1}(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\frac{D-3}{2}\right)\Gamma\left(\frac{5-D}{2}\right)}{\pi^{\frac{D-1}{2}}\Gamma\left(\frac{5-D}{2}-\ell\right)}c_{[j|i_{1}]\cdots i_{\ell}}\partial_{a}\delta^{j}_{[c}\partial_{b]}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}$ (3.67)
and simplify the tensor structure using [^29] Note that in deriving (3.2.1) we assume that $\partial_{b}$ is a derivative with respect to coordinates on the (D2)(D-2)-sphere—which we make explicit in (3.70). This allows us to factor out $1/|\vec{x}|^{2\ell+D-3}$ in (3.2.1).
$c_{[j|i_{1}]\cdots i_{\ell}}$ $\delta^{j}_{[c}\partial_{b]}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)=\frac{1}{4}\left(\partial_{b}c_{c|A_{\ell}}-\partial_{c}c_{b|A_{\ell}}-\partial_{b}c_{i_{\ell}|cA_{\ell-1}}+\partial_{c}c_{i_{\ell}|bA_{\ell-1}}\right)x^{i_{1}}\cdots x^{i_{\ell}}$
$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\partial_{c}\left(x^{i}c_{i|A_{\ell}}\frac{1}{\ell}\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})\right)-(c\leftrightarrow b)\right]$
$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\frac{1}{\ell}\partial_{c}\partial_{b}\left(x^{i}c_{i|A_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}\right)-\frac{1}{\ell}\partial_{c}\left(c_{b|A_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}\right)-(c\leftrightarrow b)\right]$
$=\frac{1}{4}\left[\partial_{b}(x^{i_{1}}\cdots x^{i_{\ell}})c_{c|A_{\ell}}+\frac{1}{\ell}\partial_{b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c|A_{\ell}}-(c\leftrightarrow b)\right]$
$=\frac{\ell+1}{2\ell}\partial_{[b}\left(x^{i_{1}}\cdots x^{i_{\ell}}\right)c_{c]|A_{\ell}}.$ (3.68)
Considering the $B_{rAB}$ component of (3.66), where $AB$ denote angular coordinates, and writing the tidal field in spherical coordinates as
@@
$\langle B_{+\,rAB}(x)\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}\frac{\ell+1}{2\ell}\frac{(\ell+1)!}{\ell}K^{(B)}_{\ell}(\omega)\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}r^{2}\partial_{r}\left(r^{-2}\frac{c_{\text{ext}}}{r^{\ell+D-4}}\right).$ (3.70)
@@
$\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rAB}$ $=\frac{2}{M_{\rm Pl}}\left[\nabla_{r}\nabla_{[i}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j]0}-\partial_{t}\nabla_{[i}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j]r}\right]$ (3.71)
$=2\operatorname{e}^{-i\omega t}\sum_{\ell=2}^{\infty}c_{\text{ext}}r^{2}\partial_{r}\big(r^{\ell-1}\big)\big(1+O(\omega r)\big)\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$.
The $O(\omega r)$ term in $\mkern 2.5mu\overline{\mkern-2.5muB\mkern-2.5mu}\mkern 2.5mu_{rAB}$ represents a far-zone, frequency-dependent correction to the static tidal field. It exhibits a different scaling in $r$ compared to the standard $r^{\ell}$ and $r^{-\ell-1}$ falloffs. Therefore, as long as we are concerned with matching the EFT in the intermediate zone, we can consistently neglect this contribution. Putting everything together and replacing $K_{\ell}^{(B)}$ by $\mu^{2\varepsilon}K_{\ell}^{(B)}$ (as we did in the scalar section), we obtain
@@
$\times\left(r^{\ell}+\mu^{2\varepsilon}\frac{K^{(B)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}\frac{\ell+1}{2\ell}\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-2}\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell+3-D}\right),$
@@
where $q_{\ell}\equiv 1/\sqrt{2(\ell-1)(D-2+\ell)}$. The components (3.75) can be used to evaluate the Weyl tensor in Regge–Wheeler gauge. In particular, we find $B_{rAB}$ to be
$B_{rAB}^{\text{RW}}$ $=\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}\left[r^{2}\partial_{r}\left(r^{-2}h_{0}(t,r)\right)-\partial_{0}h_{1}(t,r)\right]\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$ (3.76)
$=-\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}\frac{q_{\ell}\operatorname{e}^{-i\omega t}}{4}\mu^{\varepsilon}r^{1-\frac{D}{2}}\left[\left(2D-D^{2}+4r^{2}\omega^{2}\right)\Psi_{\text{RW}}(r)+4r^{2}\Psi_{\text{RW}}^{\prime\prime}(r)\right]\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$.
@@
$B_{rAB}^{\text{RW}}=-\frac{1}{M_{\rm Pl}}\sum_{\ell=2}^{\infty}q_{\ell}\operatorname{e}^{-i\omega t}\left[(\ell-1)(\ell+2-2\varepsilon)+r^{2}\omega^{2}\right]\left(B_{\text{reg}}r^{\ell}+\frac{\mu^{2\varepsilon}B_{\text{irr}}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right)\nabla_{[A}Y_{B]}^{(T)}{}_{\ell}^{m}$. (3.77)
Since the linearized Weyl tensor is gauge invariant in flat space, we can directly compare $B_{rAB}^{\text{RW}}$ with (3.72), obtained in de Donder gauge. By matching the $r^{\ell}$ and $r^{-\ell-1+2\varepsilon}$ coefficients in the two expressions, we can identify
$\frac{\ell+1}{2\ell}\frac{(\ell+1)!(D+\ell-2)}{\ell(\ell-1)}\frac{2^{\ell-1}\Gamma\left(\ell+\frac{D-1}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{K^{(B)}_{\ell}(\omega)}{M_{\rm Pl}^{2}}=\frac{B_{\text{irr}}}{B_{\text{reg}}}.$ (3.78)
@@
$\frac{3}{2}\frac{45}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=2}(\omega)=i\omega r_{s}^{6}+\omega^{2}r_{s}^{7}\left(\frac{787}{2520}-\log(\mu r_{s})\right).$ (3.87)
@@
$\frac{4}{3}\frac{18900}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=3}(\omega)$ $=i\omega r_{s}^{8}+\omega^{2}r_{s}^{9}\left(\frac{1\,727}{2\,520}-\log(\mu r_{s})\right),$ (3.88) $\frac{5}{4}\frac{11113200}{2\pi M_{\rm Pl}^{2}}K^{(B)}_{\ell=4}(\omega)$ $=i\omega r_{s}^{10}+\omega^{2}r_{s}^{11}\left(\frac{237\,529}{277\,200}-\log(\mu r_{s})\right).$
@@ 3.2.2 Even sector
As in the magnetic sector, we can expand the response kernel in fourier space similarly to (3.53),
@@
$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i\omega(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{E}+i{\omega}r_{s}\lambda_{1,\ell}^{E}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{E}+\cdots\right],$
where $\lambda_{0,\ell}^{E}$ are the static Love numbers, $\lambda_{1,\ell}^{E}$ capture the linear-in-frequency dissipative response of the object, and $\lambda_{2,\ell}^{E}$ are the dynamical Love numbers at order $O(\omega^{2})$. Since we are interested in computing $\lambda_{2,\ell}^{E}$ for Schwarzschild black holes in four-dimensional general relativity, we will set $\lambda_{0,\ell}^{E}=0$ in the following. We will allow us to consistently neglect $O(\omega^{2})$ terms in the expressions of $\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}$ and $\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle$ in (3.92).
@@
$E_{ab}$ $=C_{0a0b}=\frac{2}{M_{\rm Pl}}\left[\partial_{a}\partial_{[0}h_{b]0}-\partial_{0}\partial_{[0}h_{b]a}\right]=\frac{1}{M_{\rm Pl}}\left[-\partial_{a}\partial_{b}h_{00}+2\partial_{0}\partial_{(a}h_{b)0}+O(\omega^{2})\right],$ (3.95)
$E_{A_{\ell}}$ $=\frac{1}{M_{\rm Pl}}\left[-\partial_{A_{\ell}}h_{00}+2\partial_{0}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0}+O(\omega^{2})\right].$
@@
$E_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}}=\frac{1}{M_{\rm Pl}^{2}}\left[\partial_{A_{\ell}}h_{00}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}-2\partial_{A_{\ell}}h_{00}\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{i_{\ell}0}-2\partial_{0}\partial_{A_{\ell-1}}h_{i_{\ell}0}\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{00}+O(\omega^{2})\right].$ (3.96)
@@
$\langle E_{+\,ab}(t,\vec{x})E_{-\,A_{\ell}}\left(\tau_{2}\right)\rangle\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}}\left(\tau_{1}\right)=-\frac{1}{M_{\rm Pl}^{3}}\langle\partial_{a}\partial_{b}h_{+\,00}\partial_{A_{\ell}}h_{-\,00}\rangle\left(\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{00}-2\partial^{0}\partial^{A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{i_{\ell}0}\right)+O(\omega^{2}),$ (3.97)
@@
$\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}(t,\vec{x})=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{\rm ext}Y_{\ell}^{m}r^{\ell}=M_{\rm Pl}\operatorname{e}^{-i\omega t}\sum_{\ell}c_{j_{1}\cdots j_{\ell}}x^{j_{1}}\cdots x^{j_{\ell}}+O(\omega^{2}),$ (3.99)
@@
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\frac{D-3}{D-2}\sum_{\ell=2}^{\infty}\ell!(-i)^{\ell}K_{\ell}^{(E)}(\omega)c^{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\int\frac{{\rm d}^{D-1}\vec{p}}{(2\pi)^{D-1}}\operatorname{e}^{i\vec{p}\cdot\vec{x}}\frac{p_{i_{1}}\cdots p_{i_{\ell}}}{\vec{p}^{2}}.$ (3.100)
@@
$\langle E_{+\,ab}(t,\vec{x})\rangle_{\text{in-in}}=-\frac{\operatorname{e}^{-i\omega t}}{M_{\rm Pl}^{2}}\sum_{\ell=2}^{\infty}K_{\ell}^{(E)}(\omega)\frac{2^{\ell-2}\ell!\,\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}\frac{D-3}{D-2}c_{i_{1}\cdots i_{\ell}}\partial_{a}\partial_{b}\frac{x^{i_{1}}\cdots x^{i_{\ell}}}{|\vec{x}|^{2\ell+D-3}}.$ (3.101)
@@
$\mkern 2.5mu\overline{\mkern-2.5muE\mkern-2.5mu}\mkern 2.5mu_{rr}$ $=\frac{1}{M_{\rm Pl}^{2}}\left(-\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}+2\partial_{0}\partial_{r}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0r}\right)+O(\omega^{2})$ (3.102)
$=-\frac{1}{M_{\rm Pl}^{2}}\partial_{r}^{2}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{00}^{\text{RW}}+O(\omega^{2})=-\operatorname{e}^{-i\omega\tau}\sum_{\ell=2}^{\infty}c_{\rm ext}\,\ell(\ell-1)r^{\ell-2}Y_{\ell}^{m}+O(\omega^{2}).$